# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v2

@@ IV.2 Floquet Chern number
IV.3 Floquet energy spectrum
Figure 5: Energy spectra for (a1,b1) the Floquet Weyl-like model and (a2,b2) the Floquet Weyl semimetal with $k_{x}\!=\!k_{y}\!=\!0$. (a1,a2) $A_{0}\!=\!0.1$ nm-1; (b1,b2) $A_{0}\!=\!0.9$ nm-1. The other parameters are the same as those in Fig. 3.
The energy spectra for the Floquet Weyl-like model and the Floquet Weyl semimetal, calculated at $k_{x}=k_{y}=0$, are shown in Fig. 5. The energy gapless points, corresponding to the Weyl nodes, exhibit a dependence on the light intensity.
IV.4 Floquet Fermi arc
Figure 6: Energy spectra (solid curves) and Fermi arcs (horizontal straight line for $E\!=\!0$) for (a1,b1) the Floquet Weyl-like model and (a2,b2) the Floquet Weyl semimetal under OBCs along the $y$ direction with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. (a1,a2) $A_{0}\!=\!0.1$ nm-1; (b1,b2) $A_{0}\!=\!0.9$ nm-1. The other parameters are the same as those in Fig. 3.
@@ VI Summary
The Weyl-semimetal-like phase we propose shares key features with traditional Weyl semimetals, such as linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arcs that connect these points. Importantly, these topological behaviors manifest in real space, rather than in momentum space.
Additionally, our investigation explores the effects of high-frequency periodic driving, where we calculate Floquet Weyl nodes, layer Floquet Chern numbers, Floquet energy spectra, and Floquet Fermi arcs. We demonstrate that the Weyl-semimetal-like nodes can be tuned via high-frequency laser driving, providing a means to manipulate these topological features. This work opens new possibilities for exploring and controlling topological phases through external stimuli.
@@ Acknowledgements.
We acknowledge helpful discussions with Ching Hua Lee. R.C. acknowledges the support from the National Natural Science Foundation of China (Grant No. 12304195) and the Chutian Scholars Program in Hubei Province.

### v3

@@ V.2 Floquet Chern number
V.3 Floquet energy spectrum
Figure 9: Energy spectra for (a1, b1) the Floquet Weyl-like model and (a2, b2) the Floquet Weyl semimetal with $k_{x}=k_{y}=0$. Panels (a1, a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1, b2) to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines mark the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 6.
The energy spectra for the Floquet Weyl-like model and the Floquet Weyl semimetal, calculated at $k_{x}\!=\!k_{y}\!=\!0$, are shown in Fig. 9. The two green vertical dashed lines are used to label the positions of the two Weyl-like (or Weyl) points. The energy gapless points, corresponding to the Weyl nodes, exhibit a dependence on the light intensity.
V.4 Floquet Fermi arc
Figure 10: Energy spectra (solid curves) and Fermi arcs (horizontal line at $E\!=\!0$) for (a1, b1) the Floquet Weyl-like model and (a2, b2) the Floquet Weyl semimetal under open boundary conditions (OBCs) along the $y$ direction, with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. Panels (a1, a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1, b2) to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines indicate the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 6.
@@ VII Summary
The Weyl-semimetal-like phase we propose shares key features with traditional Weyl semimetals, such as linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arc that connect these points. Importantly, these topological behaviors manifest in real space, rather than in momentum space.
Moreover, we investigate the effects of interlayer coupling by calculating the local density of states, the layer Hall conductance, and the total 3D Hall conductivity. Our results show that the Weyl-semimetal-like phase remains robust under weak and moderate interlayer couplings.
Additionally, our investigation explores the effects of high-frequency periodic driving, where we calculate Floquet Weyl nodes, Floquet Chern numbers, Floquet energy spectra, and Floquet Fermi arcs. We demonstrate that the Weyl-semimetal-like nodes can be tuned via high-frequency laser driving, providing a means to manipulate these topological features. This work opens new possibilities for exploring and controlling topological phases through external stimuli.
@@ Acknowledgements.
We acknowledge helpful discussions with Ching Hua Lee. F.Q. acknowledges the support from the Jiangsu Specially-Appointed Professor Program in Jiangsu Province and the Doctoral Research Start-Up Fund of Jiangsu University of Science and Technology. R.C. acknowledges the support from the National Natural Science Foundation of China (Grant No. 12304195) and the Chutian Scholars Program in Hubei Province.